# Mixed-state geometric phases of coherent and squeezed spin states

### v2

@@ II.4 Interferometric geometric phase
There is another approach for constructing a geometric phase for mixed states. Inspired by optical processes in the Mach-Zehnder interferometer, Ref. Sjöqvist et al. (2000) assigned a phase to a mixed quantum state represented by the density matrix $\rho$ undergoing a unitary evolution $U(t)$. Explicitly, if a system evolves according to $\rho(t)=U(t)\rho(0)U^{\dagger}(t)$, it gives rise to a total phase
@@
Notably, the transformation from $\rho(t)$ to $\rho(t+\mathrm{d}t)$ is $\rho(t+\mathrm{d}t)=U(t+\mathrm{d}t)\rho(0)U^{\dagger}(t+\mathrm{d}t)=U(t+\mathrm{d}t)U^{\dagger}(t)\rho(t)U(t)U^{\dagger}(t+\mathrm{d}t)$. Thus, by definition, the relative phase between the mixed states at $t$ and $t+\mathrm{d}t$ is $\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]$. In this framework Sjöqvist et al. (2000), a parallel-transport condition of $U(t)$ is proposed as follows. $\rho(t+\mathrm{d}t)$ is required to be instantaneously “in phase” with $\rho(t)$, which means $\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]=0$. In the differential form, the condition yields
@@
To specify parallel-transport by the evolution according to $U(t)$ more deterministically, the condition (18) has been strengthened as Sjöqvist et al. (2000)
@@
where $|n(t)\rangle$ is the $n$th eigen-vector of $\rho(t)$. We emphasize that the parallel-transport condition (18) for the IGP fundamentally differs from Uhlmann’s condition in Eq. (13), due to the U$(N)$ gauge dependence of the phase factor $\mathcal{U}$. Crucially, introducing $\mathcal{U}$ via $W=\sqrt{\rho}\mathcal{U}$ can violate the parallel-transport condition (18) Hou et al. (2023b).
Under the parallel-transport condition, the total phase accumulated during the evolution is the geometric phase referred to as the IGP:
$\theta_{\text{G}}(t)=\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(0)U(t)\right].$ (20)
@@
As a consequence, the total phase contains only the geometric contribution from the IGP. We note that the strengthened condition (19) will result in vanishing integrand in the above expression, thereby resulting the same vanishing dynamic phase.
In certain situations, it may be difficult for unitary evolutions to fully satisfy the parallel-transport condition (18). Nevertheless, the above discussion can be extended to general unitary evolution beyond parallel transport, which results in non-adiabatic phases Zhang et al. (2023). For pure states, a similar formalism is known as the quantum kinematic approach to the geometric phase Mukunda and Simon (1993). Recently, a similar non-adiabatic geometric phase for mixed states has been discussed in Ref. Mohammadi Almas and Najarbashi (2024). We follow those ideas and define the non-adiabatic IGP for any unitary process by subtracting the non-adiabatic generalization of the dynamic phase from the total phase. Explicitly,
$\theta_{\text{G}}^{\prime}(t)=\theta_{\text{T}}(t)-\theta_{\text{NG}}(t).$ (22)
Here
$\theta_{\text{NG}}(t)=-\mathrm{i}\int_{0}^{t}\mathrm{d}t^{\prime}\text{Tr}\left[\rho(t^{\prime})\dot{U}(t^{\prime})U^{\dagger}(t^{\prime})\right]$ (23)
is considered as a non-geometric phase because if $U(t)$ represents a dynamic evolution, $\theta_{\text{NG}}(t)$ reduces to the dynamic phase. Moreover, if $U(t)$ satisfies the parallel-transport condition, $\theta_{\text{NG}}(t)=0$, and the total phase equals to the geometric phase given by the IGP.
@@ II.5 Difference between the two geometric phases
We emphasize that the two aforementioned geometric phases for mixed states are physically and mathematically distinct. Although the theory of IGP can also be formulated equivalently using a purification formalism of density matrices Sjöqvist et al. (2000), any phase factor $\mathcal{U}$ is ruled out in the derivation. Specifically, only the trivial decomposition $W=\sqrt{\rho}$ is permissible for the IGP because otherwise the condition (18) would be violated.
In contrast, the generation of the Uhlmann phase inherently relies on the phase factor $\mathcal{U}$ according to the underlying fiber-bundle structure. All purifications of a given density matrix $\rho$ thus span a U$(N)$ fiber space. Consequently, the Uhlmann phase is naturally and directly connected to the topological properties of the U$(N)$ principal bundle, a feature absent in the IGP. Explicitly, the Uhlmann phase corresponds to the Uhlmann holonomy of the Uhlmann bundle of density matrices, similar to the Berry phase revealing the Berry holonomy of the fiber bundle of pure states Guo et al. (2020). Moreover, the Uhlmann phase requires a cyclic process but is guaranteed to be gauge invariant Sjöqvist et al. (2000); Hou et al. (2023a) while the IGP and its generalization do not impose those constraints. In the following, we will compare the differences between the Uhlmann phase and IGP through specific examples of the CSS and SSS.
III Geometrical Phases of CSS
@@ III.1 The Uhlmann phase
Using Eq. (III.1), the Uhlmann phase generated in a cyclic process can be evaluated. The $A_{U}$ of the CSS only differs by a minus sign from that of the spin-$j$ system discussed in Ref. Hou et al. (2021b). Therefore, their Uhlmann phases only differ by a minus sign. As explained in Sec. II, this is because the CSS is related to the spin-$j$ system with a particular Hamiltonian. Since the cases with $j=\frac{1}{2}$ and $1$ of the corresponding spin-$j$ systems have already been studied in Ref. Hou et al. (2021b), we will focus here on the $j=\frac{3}{2}$ CSS. In Appendix A, we show that (I) the Uhlmann phase reduces to the Berry phase associated with the ground state as $T\rightarrow 0$, thereby confirming the correspondence explained in Ref. Wang et al. (2023), and (II) the $T\rightarrow 0$ expressions of the CSSs agree with those of Ref. Chaturvedi et al. (1987) for pure states.
@@
$\theta_{U}$ $=\arg\operatorname{Tr}\left[\frac{1}{Z}\mathrm{e}^{\beta\omega_{0}J_{x}}\mathcal{P}e^{\mathrm{i}\chi{\oint}_{\gamma}J_{z}\mathrm{d}\phi}\right],$ (44)
@@
We plot $\theta_{U}$ as a function of $T$ in Figure 1. In the low temperature limit, $\theta_{U}=\pi$, which indicates that the topological properties of the system change after a cyclic evolution along the equator. More specifically, the final purification (in terms of the fiber in the bundle language) is “antiparallel” to the initial purification. As temperature increases and crosses the first critical temperature $T^{1}_{c}\approx 0.321\omega_{0}$, the value of $\theta_{U}$ suddenly drops to 0, implying that a topological change occurs in the evolution along the equator on the parameter space. The final purification now becomes “parallel” to the initial purification. As temperature continues to rise, subsequent topological transitions occur at $T^{2}_{c}\approx 0.376\omega_{0}$ and $T^{3}_{c}\approx 0.493\omega_{0}$. In the high temperature limit, $\theta_{U}$ remains zero, meaning the associated Uhlmann holonomy is topologically trivial. This is reasonable since $\lim_{T\rightarrow 0}\rho=\frac{1}{4}1_{4\times 4}$. Therefore, the evolution loop $\rho(t)$ essentially becomes a single point with its horizontal lift $W(t)$ also reduced to a single point Hou et al. (2021b). Hence, there is no difference between the initial and final fibers anymore. We mention that bosonic and fermionic coherent states at finite temperatures, in contrast, have been shown to have smooth behavior of the Uhlmann phase without any jumps Wang et al. (2023).
@@ III.2 The interferometric geometric phase
A possible choice is $\tan^{2}\frac{\theta}{2}\mathrm{d}\phi=0$, which implies either $\theta=0$ or $\mathrm{d}\phi=0$. The former leads to $\zeta=0$, indicating that $D(\zeta)=1$, which is a trivial evolution. Therefore, we choose the latter and set $\phi$ to be constant. This represents a longitude on the unit-sphere parameter space shown by the big circle in Figure 2. We remark that the same result can be obtained if the strengthened parallel-transport condition (19) is used instead.
Figure 2: The great circle represents a longitude with constant $\phi$ on the unit-sphere parameter space, which satisfies the parallel-transport condition (III.2). In contrast, the twisting curve denotes a simple evolutionary path with $\theta=\phi$, which does not fulfill the parallel-transport condition. Both paths start from the north pole, depicted by the top blue point.
@@
In the bottom panel of Figure 3, we choose a special case with $\theta_{f}=\frac{3}{4}\pi$ to show the discontinuity of the IGP at $T_{c}=0.408\omega_{0}$. Similar to the case of the Uhlmann phase, this can be recognized as a geometric phase transition induced by temperature. The only difference is that, in general, the IGP is not directly related to the topological properties of the system. Furthermore, finite-temperature geometric phase transitions of the IGP seem to be less common than those of the Uhlmann phase. For a two-level Hermitian quantum system, a finite change in temperature in general does not induce any jump of the IGP Andersson et al. (2016); Hou et al. (2023a). For a three-level system, a careful design of the evolution process can make it happen Hou et al. (2023a). Our results of the CSS thus illustrate additional transitions of the IGP and show that the CSS is suitable for studying interesting properties of mixed-state geometric phases.
Figure 3: (Top panel) Contour plot of the IGP $\theta_{G}$ versus temperature $T$ and $\theta_{f}$ for $j=\frac{3}{2}$. The geometric phase transitions when $\theta_{f}\in(\frac{\pi}{2},\frac{3\pi}{2})$ are indicated by the boundaries separating the two values of $\theta_{G}$. (Bottom panel) $\theta_{G}$ versus temperature $T$ for the case with $\theta_{f}=\frac{3\pi}{4}$.
The above studies can also be extended to generic unitary evolutions which do not necessarily satisfy the parallel-transport condition (III.2). As previously noted, the non-adiabatic IGP is generated during these evolutions according to Eq. (22):
$\theta^{\prime}_{\text{G}}(t)=\arg\operatorname{Tr}[\rho(0)D(t)]+\mathrm{i}\int_{0}^{t}\operatorname{Tr}\left[\rho(0)D^{\dagger}(t^{\prime})\dot{D}(t^{\prime})\right]\mathrm{d}t^{\prime}.$ (48)
The evolution is characterized by a trajectory $\theta=f(\phi)$ or, conversely, $\phi=g(\theta)$ in the parameter space. To contrast with the IGP, we focus on evolution during which $\theta^{\prime}_{\text{G}}$ exhibits finite jumps. Specifically, a simple evolutionary curve $C$ defined by $\theta(t)=\phi(t)$ with $\phi\in[0,\pi]$ is selected, as shown by the twisting curve in Fig. 2. This curve starts at $\theta=0$ when $t=0$ and ends at any point $(\theta_{f},\phi_{f})$ along the path, where $\theta_{f}\equiv\theta(t_{\text{final}})$ and $\phi_{f}\equiv\phi(t_{\text{final}})$. In the top panel of Figure 4, we present the contour plot of $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\phi_{f}$ for the $j=\frac{3}{2}$ CSS. Although $\theta^{\prime}_{\text{G}}$ is not quantized, its value still undergoes quantized jumps. To better illustrate those jumps, we plot $\theta^{\prime}_{\text{G}}$ vs $T$ with $\theta_{f}=\frac{3\pi}{4}$ and $\theta^{\prime}_{\text{G}}$ vs $\theta_{f}$ with $T=0.6\omega_{0}$, respectively, in the bottom panels of Fig. 4.
Interestingly, the critical temperature $T_{c}=0.408\omega_{0}$ with a $\pi$-jump when $\theta_{f}=\frac{3\pi}{4}$ is identical to that observed along the path with constant $\phi$ which respects the parallel-transport condition, as shown in the bottom panel of Fig. 3. This can be understood by noting that in the non-geometric phase, the second term on the right-hand-side of Eq. (48) is always continuous and vanishes during parallel transport. The discontinuity can only arise from the first term, which appears in both cases with and without the parallel-transport condition. At fixed temperature $T=0.6\omega_{0}$ but with varying end points labeled by $\theta_{f}$, $\theta^{\prime}_{\text{G}}$ also experiences a $\pi$-jump at about $\theta_{f}\approx 2.05$ rad, as indicated by the right panel in the bottom of Fig. 4. Therefore, the IGP and its non-adiabatic generalization of the CSS offer interesting predictions for future experimental investigations.
Figure 4: (Top panel) The non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\phi_{f}$, where $\theta_{f}=\phi_{f}$ corresponding to $\theta_{f}\in[0,\frac{3\pi}{4}]$. The black curve indicates geometrical phase transitions. (Bottom panels) $\theta^{\prime}_{\text{G}}$ vs $T$ at $\theta_{f}=\frac{3\pi}{4}$ (left) and $\theta^{\prime}_{\text{G}}$ vs $\theta_{f}$ at $T=0.6\omega_{0}$ (right). In both two cases, $\theta^{\prime}_{\text{G}}$ exhibits $\pi$-jumps.
IV GEOMETRIC PHASES OF SSS
@@ IV.1.1 Uhlmann Phase
For the SSS shown in Eq. (5), the parameter space is one-dimensional. Therefore, we need to determine the value of $\Theta$ to define a cyclic process for density matrices, which is essential for generating the Uhlmann phase. The density matrix of a system in thermal equilibrium governed by $H(\Theta)$ show in Eq. (9) is given by $\rho(\Theta)=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta)}$ with $Z=\operatorname{Tr}(\mathrm{e}^{-\beta H(\Theta)})$. By examining Eq. (6), we observe that the quadratic term of $J_{x}$ appears in the exponent, making it difficult to derive a general expression for the SSS. Therefore, for a concise illustration of the geometric phases of the SSS, we will focus our studies on a fixed value of $j$. Since the behavior of the IGP for a two-level system ($j=\frac{1}{2}$) is relatively simple Andersson et al. (2016), we will begin with $j=1$ for our analysis.
@@
$S(\Theta)=\left(\begin{array}[]{ccc}\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}&0&-\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}\\ 0&e^{-\frac{\mathrm{i}\Theta}{2}}&0\\ -\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}&0&\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}\\ \end{array}\right).$ (63)
A straightforward calculation shows that when $\Theta=4\pi$, $S(4\pi)=\mathbf{1}_{3\times 3}$, which leads to $\rho(4\pi)=\rho(0)$ and gives a cyclic process. The $4\pi$ periodicity also contrasts the $j=1$ SSS from an ordinary spin-1 state.
@@
Nevertheless, numerical results can be obtained by using the Trotter-Suzuki approximation Trotter (1959); Suzuki (1976) to tackle the path-ordered expression. Our findings are presented in Figure 5, where $\theta_{U}$ is plotted as a function of $T$. The Uhlmann phase of SSS, similar to that of CSS, exhibits a $\pi$-valued jump at a finite temperature $T_{c}\approx 0.68\omega_{0}$, indicating the existence of a temperature-induced topological phase transition. This result reveals that at finite temperatures, the initial and final purifications no longer remain parallel, and the “angle” between them undergoes an abrupt transition from “parallel” to “antiparallel”. As shown in Fig. 5, this explains the jump in the Uhlmann phase from $\pi$ to 00 at $T_{c}$. In the infinite-temperature limit, $\theta_{U}=0$ again. This is reasonable since $\lim_{T\rightarrow+\infty}\rho$ is proportional to the identity matrix, and the associated horizontal lift shrinks to a single point, resulting in $\lim_{T\rightarrow+\infty}\theta_{U}=0\mod 2\pi$ Hou et al. (2021a).

### v3

@@ II.4 Interferometric geometric phase
There is another approach for constructing a geometric phase for mixed states. Inspired by optical processes in the Mach-Zehnder interferometer, Ref. [27] assigned a phase to a mixed quantum state represented by the density matrix $\rho$ undergoing a unitary evolution $U(t)$. Explicitly, if a system evolves according to $\rho(t)=U(t)\rho(0)U^{\dagger}(t)$, it gives rise to a total phase
@@
Notably, the transformation from $\rho(t)$ to $\rho(t+\mathrm{d}t)$ is $\rho(t+\mathrm{d}t)=U(t+\mathrm{d}t)\rho(0)U^{\dagger}(t+\mathrm{d}t)=U(t+\mathrm{d}t)U^{\dagger}(t)\rho(t)U(t)U^{\dagger}(t+\mathrm{d}t)$. Thus, by definition, the relative phase between the mixed states at $t$ and $t+\mathrm{d}t$ is $\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]$. In this framework [27], a parallel-transport condition of $U(t)$ is proposed as follows. $\rho(t+\mathrm{d}t)$ is required to be instantaneously “in phase” with $\rho(t)$, which means $\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]=0$. In the differential form, the condition yields
@@
To specify the parallel transport by the evolution according to $U(t)$ more deterministically, the condition (19) has been strengthened as [27]
@@
where $|n(t)\rangle$ is the $n$th eigen-vector of $\rho(t)$. We emphasize that the parallel-transport condition (19) or its strengthened form (20) for the IGP fundamentally differs from Uhlmann’s condition in Eq. (14), due to the U$(N)$ gauge dependence of the phase factor $\mathcal{U}$. Crucially, introducing $\mathcal{U}$ via $W=\sqrt{\rho}\mathcal{U}$ can violate the parallel-transport condition (19) [64].
In general, the IGP parallel-transport condition shown in Eq. (20) can always be satisfied, as demonstrated in Refs. [27, 32]. For instance, in our earlier discussion of the one-axis SSS, we introduced the operator $\tilde{S}$. It can be shown that for all $n$,
$\langle n(t)|\dot{\tilde{S}}(t)\tilde{S}^{\dagger}(t)|n(t)\rangle=\langle n(t)|\sum_{m}\frac{\mathrm{d}|m(t)\rangle}{\mathrm{d}t}\langle m(t)|n(t)\rangle$
-- $\langle n(t)|\sum_{m}\langle m(t)|\frac{\mathrm{d}}{\mathrm{d}t}|m(t)\rangle|n(t)\rangle\delta_{mn}=0.$ (21)
Hence, the IGP parallel-transport condition is automatically satisfied, regardless of the explicit form of $S$.
Under the parallel-transport condition, the phase accumulated during the evolution is the geometric phase referred to as the IGP:
$\theta_{\text{G}}(t)=\arg\text{Tr}\left[\rho(0)U(t)\right].$ (22)
@@
As a consequence, the total phase contains only the geometric contribution from the IGP. We note that the strengthened condition (20) will result in vanishing integrand in the above expression, thereby resulting the same vanishing dynamic phase. We mention that general unitary evolution beyond parallel transport could result in non-adiabatic phases [65]. There is a formalism known as the quantum kinematic approach to the geometric phase for pure states [66], and a similar non-adiabatic geometric phase for mixed states has been discussed in Ref. [67].
@@ II.5 Difference between the two geometric phases
We emphasize that the two aforementioned geometric phases for mixed states are physically and mathematically distinct. Although the theory of IGP can also be formulated equivalently using a purification formalism of density matrices [27], any phase factor $\mathcal{U}$ is ruled out in the derivation. Specifically, only the trivial decomposition $W=\sqrt{\rho}$ is permissible for the IGP because otherwise the condition (19) would be violated.
In contrast, the generation of the Uhlmann phase inherently relies on the phase factor $\mathcal{U}$ according to the underlying fiber-bundle structure. All purifications of a given density matrix $\rho$ thus span a U$(N)$ fiber space. Consequently, the Uhlmann phase is naturally and directly connected to the topological properties of the U$(N)$ principal bundle, a feature absent in the IGP. Explicitly, the Uhlmann phase corresponds to the Uhlmann holonomy of the Uhlmann bundle of density matrices, similar to the Berry phase revealing the Berry holonomy of the fiber bundle of pure states [62]. Moreover, the Uhlmann phase requires a cyclic process but is guaranteed to be gauge invariant [27, 38] while the IGP and its generalization do not impose those constraints. In the following, we will compare the differences between the Uhlmann phase and IGP through specific examples of the CSS and SSS.
III Mixed-state Geometrical Phases of CSS
@@ III.1 The Uhlmann phase
Using Eq. (III.1), the Uhlmann phase generated in a cyclic process can be evaluated. The $A_{U}$ of the CSS only differs by a minus sign from that of the spin-$j$ system discussed in Ref. [68]. Therefore, their Uhlmann phases only differ by a minus sign. As explained in Sec. II, this is because the CSS is related to the spin-$j$ system with a particular Hamiltonian. Since the cases with $j=\frac{1}{2}$ and $1$ of the corresponding spin-$j$ systems have already been studied in Ref. [68], we will focus here on the $j=\frac{3}{2}$ CSS. In Appendix A, we show that (I) the Uhlmann phase reduces to the Berry phase associated with the ground state as $T\rightarrow 0$, thereby confirming the correspondence explained in Ref. [21], and (II) the $T\rightarrow 0$ expressions of the CSSs agree with those of Ref. [57] for pure states.
@@
$\theta_{U}$ $=\arg\operatorname{Tr}\left[\frac{1}{Z}\mathrm{e}^{\beta\omega_{0}J_{x}}\mathcal{P}\mathrm{e}^{\mathrm{i}\chi{\oint}_{\gamma}J_{z}\mathrm{d}\phi}\right],$ (44)
@@
We plot $\theta_{U}$ as a function of $T$ in Figure 1. In the low temperature limit, $\theta_{U}=\pi$, which indicates that the topological properties of the system change after a cyclic evolution along the equator. More specifically, the final purification (in terms of the fiber in the bundle language) is “antiparallel” to the initial purification. As temperature increases and crosses the first critical temperature $T^{1}_{c}\approx 0.321\omega_{0}$, the value of $\theta_{U}$ suddenly drops to 0, implying that a topological change occurs in the evolution along the equator on the parameter space. The final purification now becomes “parallel” to the initial purification. As temperature continues to rise, subsequent topological transitions occur at $T^{2}_{c}\approx 0.376\omega_{0}$ and $T^{3}_{c}\approx 0.493\omega_{0}$. In the high temperature limit, $\theta_{U}$ remains zero, meaning the associated Uhlmann holonomy is topologically trivial. This is reasonable since $\lim_{T\rightarrow 0}\rho=\frac{1}{4}1_{4\times 4}$. Therefore, the evolution loop $\rho(t)$ essentially becomes a single point with its horizontal lift $W(t)$ also reduced to a single point [68]. Hence, there is no difference between the initial and final fibers anymore. We mention that bosonic and fermionic coherent states at finite temperatures, in contrast, have been shown to have smooth behavior of the Uhlmann phase without any jumps [21].
@@ III.2 The interferometric geometric phase
A possible choice is $\tan^{2}\frac{\theta}{2}\mathrm{d}\phi=0$, which implies either $\theta=0$ or $\mathrm{d}\phi=0$. The former leads to $\zeta=0$, indicating that $D(\zeta)=1$, which is a trivial evolution. Therefore, we choose the latter and set $\phi$ to be constant, which corresponds to a longitude on the unit-sphere parameter space. We remark that the strengthened parallel-transport condition (20) is also satisfied along the path.
@@
In the bottom panel of Figure 2, we choose a special case with $\theta_{f}=\frac{3}{4}\pi$ to show the discontinuity of the IGP at $T_{c}=0.408\omega_{0}$. Similar to the case of the Uhlmann phase, this can be recognized as a geometric phase transition induced by temperature. The only difference is that, in general, the IGP is not directly related to the topological properties of the system. Furthermore, finite-temperature geometric phase transitions of the IGP seem to be less common than those of the Uhlmann phase. For a two-level Hermitian quantum system, a finite change in temperature in general does not induce any jump of the IGP [39, 38]. For a three-level system, a careful design of the evolution process can make it happen [38]. Our results of the CSS thus illustrate additional transitions of the IGP and show that the CSS is suitable for studying interesting properties of mixed-state geometric phases.
Figure 2: (Top panel) Contour plot of the IGP $\theta_{G}$ versus temperature $T$ and $\theta_{f}$ for $j=\frac{3}{2}$. The geometric phase transitions when $\theta_{f}\in(\frac{\pi}{2},\frac{3\pi}{2})$ are indicated by the boundaries separating the two values of $\theta_{G}$. (Bottom panel) $\theta_{G}$ versus $T$ for the case with $\theta_{f}=\frac{3\pi}{4}$.
IV Mixed-state GEOMETRIC PHASES OF SSS
@@ IV.1.1 Uhlmann Phase
For the SSS shown in Eq. (5), the parameter space is one-dimensional. Therefore, we need to determine the value of $\Theta$ to define a cyclic process for density matrices, which is essential for generating the Uhlmann phase. The density matrix of a system in thermal equilibrium governed by $H(\Theta)$ show in Eq. (10) is given by $\rho(\Theta)=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta)}$ with $Z=\operatorname{Tr}(\mathrm{e}^{-\beta H(\Theta)})$. By examining Eq. (6), we observe that the quadratic term of $J_{x}$ appears in the exponent, making it difficult to derive a general expression for the SSS. Therefore, for a concise illustration of the geometric phases of the SSS, we will focus our studies on a fixed value of $j$. Since the behavior of the IGP for a two-level system ($j=\frac{1}{2}$) is relatively simple [39], we will begin with $j=1$ for our analysis. Crucially, Eq. (16) indicates that the Uhlmann connection $A_{U}$ only depends on $\rho(t)$ and its eigenvalues and eigenstates. Given that $\rho(t)=S(\Theta(t))\rho(0)S^{\dagger}(\Theta(t))=\tilde{S}(\Theta(t))\rho(0)\tilde{S}^{\dagger}(\Theta(t))$, the expressions following $S(\Theta)$ must be identical to those following $\tilde{S}(\Theta)$. Consequently, the Uhlmann phase calculated above is the same as that from the evolution governed by $\tilde{S}(\Theta)$. The identification allows for a fair comparison with the IGP discussed later. We mention that the equivalent evaluation via $S(\Theta)$ simplifies the calculation without the distraction from the additional pure-state geometric phase associated with each $|jm\rangle$, which does not contribute to the final expression of the Uhlmann phase.
@@
$S(\Theta)=\left(\begin{array}[]{ccc}\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta}{2}}&0&-\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta}{2}}\\ 0&\mathrm{e}^{-\frac{\mathrm{i}\Theta}{2}}&0\\ -\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta}{2}}&0&\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta}{2}}\\ \end{array}\right).$ (62)
A straightforward calculation shows that when $\Theta=4\pi$, $S(4\pi)=\mathbf{1}_{3\times 3}$, which leads to $\rho(4\pi)=\rho(0)$ and gives a cyclic process. The $4\pi$ periodicity also contrasts the $j=1$ SSS from an ordinary spin-1 state. Hereafter, we consider $\Theta\in[0,4\pi]$ for the $j=1$ SSS.
@@
Nevertheless, numerical results can be obtained by using the Trotter-Suzuki approximation [69, 70] to tackle the path-ordered expression. Our findings are presented in Figure 3, where $\theta_{U}$ is plotted as a function of $T$. The Uhlmann phase of SSS, similar to that of CSS, exhibits a $\pi$-valued jump at a finite temperature $T_{c}\approx 0.68\omega_{0}$, indicating the existence of a temperature-induced topological phase transition. This result reveals that at finite temperatures, the initial and final purifications no longer remain parallel, and the “angle” between them undergoes an abrupt transition from “parallel” to “antiparallel”. As shown in Fig. 3, this explains the jump in the Uhlmann phase from $\pi$ to 00 at $T_{c}$. In the infinite-temperature limit, $\theta_{U}=0$ again. This is reasonable since $\lim_{T\rightarrow+\infty}\rho$ is proportional to the identity matrix, and the associated horizontal lift shrinks to a single point, resulting in $\lim_{T\rightarrow+\infty}\theta_{U}=0\mod 2\pi$ [24].